OpenStudy (anonymous):

What is the solution to this equation? |2y-5|=7

OpenStudy (anonymous):

write 2 equations. \[-2y+5=7 \] and \[2y-5=7\]. from there, it's basic algebra

OpenStudy (anonymous):

-2y+5=7, Y=-1.5 2y-5=7, Y=6

OpenStudy (anonymous):

-1 or 6

OpenStudy (anonymous):

eeyup, absolute value equations should always end with 2 solutions